# Mixing, segregation, and collapse transitions of interacting copolymer rings

### v1

@@ 1 Introduction
In this paper we investigate a model of two ring polymers that are required to be close together in space. The monomers in the two rings are different and this can lead to a mixed phase where the two rings interpenetrate [JOTW22]. In addition, either or both of the rings can collapse to form a compact ball. This model is connected to the behaviour of a diblock copolymer.
The synthesis of catenated ring polymers [Ohta2008, Cao2015catenated], often in $AB$-diblock form [niu2009polycatenanes] with one ring the $A$-block and the other ring the $B$-block, presents a novel situation with a phase diagram admitting several possible phases, including segregated, collapsed and mixed phases in a dilute solution, in addition to additional phases when the diblock catenane is in a melt or in confinement [Cao2015catenated]. The linking of components in a catenane mix results in interlocking ring polymer sheets [Luengo2024shape] resembling chainmail and held together by mechanically bonded rings due to the physical linking of the component ring polymers. Numerical simulations of catenated ring copolymers show that the rings are pulled together by catenation, giving rise to smaller radii of gyration (when compared to a single ring of the same length) and reduced distance between the centres of mass [SC14, JOTW22]. Furthermore, it has been shown that ring-shaped polycatenanes have the capacity to retain a relevant quantity of twist, which significantly alters their metric and local properties [tubiana2022circular].
@@
A cubic lattice model of two ring copolymers in a dilute solution that may be catenated consists of two interacting ring polymers (namely, lattice polygons) as illustrated in figure 1. The two polygons are held in proximity, and we examine metric and linking properties using Monte Carlo simulations. This approach is broadly similar to the study in reference [SC14], but with additional (repulsive or attractive) short-ranged forces between monomers in the two respective component polygons, which are also (independently) self-interacting if there are forces between monomers in each. This distinguishes the two polygons, so that, when catenated, they are a model of an $AB$-diblock ring polymer, with one ring the $A$-block and the other ring the $B$-block.
If the self-interacting forces between monomers in each component polygon are large, then they go through $\theta$-points to collapsed regimes, while, if the forces between monomers in each block (one in the $A$-block, and the other in the $B$-block) are large, then the two component polygons should interpenetrate each other, moving into a mixed regime. If either of the $A$- or $B$-blocks are collapsed, then they are in a collapsed or globular phase with increased self-entanglements as observed by the increased probability that they are non-trivial knots [tesi1994knotting]. In the mixed regime the probability that the components form a catenane is increased [orlandini2004entangled, panagiotou2011study, JOTW22]. We investigate this by examining the linking probability of the $A$- and $B$-blocks as a function of the interaction between the components.
We implement the model in figure 1 by considering two lattice polygons (or closed self-avoiding walks) of the same length on the simple cubic lattice. The two polygons are held together in the lattice by having at least one vertex in one polygon (this is the $A$-block) a unit distance from a vertex in the other polygon (namely the $B$-block). There are three energy terms in this model. If two monomers of type $A$ are unit distance apart and not connected by an edge of the polygon, then they form an $AA$-contact and they contribute an energy $\epsilon_{AA}$ to the total energy of the system. Similarly, two monomers of type $B$ a unit distance apart form a $BB$-contact and contribute an energy $\epsilon_{BB}$, and a pair of monomers of types $A$ and $B$ a unit distance apart form an $AB$-contact and contribute an energy $\epsilon_{AB}=\epsilon_{m}$. If $\epsilon_{AA}<0$ this will favour collapse of the $A$-ring into a globular phase, and similarly, $\epsilon_{BB}<0$ will favour collapse of the $B$-ring. If $\epsilon_{AB}<0$, this will favour intermingling of the two polygons and a transition to a mixed phase where we examine the linking probability of the two rings [JOTW22].
@@
If we set $\epsilon_{AA}=\epsilon_{BB}$, then both rings will collapse at the same temperature. In that case we expect at least three phases: A segregated-expanded phase where the rings are largely separated in space and each is expanded as in a good solvent, a segregated-collapsed phase where the two rings are largely separated but each ring has collapsed to a compact ball, and a mixed phase where the two rings interpenetrate to a significant extent. The phase diagram is illustrated in figure 2. We expect three phase boundaries in the phase diagram. The phase transition between the segregated-expanded and segregated-collapsed phases when $\epsilon_{AB}=0$ is relatively well understood. It is the collapse transition for a ring polymer. It is thought to be second order and there are good numerical estimates of its location [bennett1998exact, tesi1996interacting]. The estimates are
@@
[…]
We use several techniques to investigate this phase diagram. In section 3, we use rigorous arguments to establish the existence of a phase boundary as $\beta_{m}$ is increased, for all $\beta_{c}<\infty$. We write this phase boundary as $\beta_{m}=\beta_{m}^{*}(\beta_{c})$. In the same section, we show that, if a collapse transition exists as $\beta_{c}$ is increased with $\beta_{m}=0$ (say at $\beta_{c}=\beta_{c}^{*}$), then a transition occurs at the same value of $\beta_{c}$ for all $\beta_{m}<\beta_{m}(\beta_{c}^{*})$.
Our rigorous results leave many questions unanswered and we use Monte Carlo methods to obtain additional information. There are several quantities that one can examine to probe the details of the phase diagram. These include thermodynamic quantities such as the energy and heat capacity, and metric properties such as the radii of gyration of the rings (to detect collapse), and the mean distance between the centres of gravity of the rings (to detect mixing). In addition, we can probe the extent of intermingling or mutual entanglement by calculating the linking probability of the two rings. We proceed by sampling and collecting data on the model along lines crossing the phase boundaries in the phase diagram, using a multiple Markov Chain implementation [geyer1995annealing, tesi1996monte] of the Metropolis algorithm [metropolis1953monte] using Verdier-Stockmayer [verdier1962monte] and pivot [madras1990monte] elementary moves which can change the knot type of lattice polygons and linking complexity between lattice polygons.
@@ 2 The model
Mutual interactions between the two polygons are mutual contacts (pairs of vertices $(v_{A}\in A,v_{B}\in B)$ such that the distance between $v_{A}$ and $v_{B}$ is one lattice unit). As above, there is an interaction energy $\epsilon_{m}$ for each mutual contact, and the total interaction energy associated with a given configuration having $k_{m}$ mutual contacts is $E_{m}=\epsilon_{m}k_{m}$. In addition, each block is also subject to an effective self-attraction due to the quality of the solvent. By defining a contact within block A as a pair of non-consecutive vertices that are a unit distance apart, we have $E_{A}=\epsilon_{A}k_{A}$ where $\epsilon_{A}<0$ is an effective attractive energy associated with each contact within block $A$ and $k_{A}$ is the number of such contacts. Similarly, for block $B$ we have $E_{B}=\epsilon_{B}k_{B}$ with $\epsilon_{B}<0$.
@@
In equation (2) the function $p_{2n}^{(2)}(k_{m},k_{A},k_{B})$ denotes the number of configurations of the diblock polygon with total length $2n$ having $k_{m}$ mutual contacts, and $k_{A}$, $k_{B}$ (self-)contacts within blocks $A$ and $B$ respectively. This general form accounts for the possibility of the two blocks having different affinities with the solvent. If we restrict to the simpler case of $\epsilon_{A}=\epsilon_{B}=\epsilon_{c}$ and define $k_{c}$ to be the total number of contacts, i.e. $k_{c}=k_{A}+k_{B}$, then the partition function (2) simplifies to
@@ 2.1 Preliminary considerations
Above a critical value $\beta_{m}^{*}$, the system is expected to be in a Mixed (M) phase where the two rings are significantly intermingled (the density of mutual contacts $\langle k_{m}\rangle/2n$ is now positive in the thermodynamic limit). The value of $\beta_{m}^{*}$ is $0.31$ for $\beta_{c}=0$  [JOTW22] but, for $\beta_{c}\neq 0$, we expect $\beta_{m}^{*}=\beta_{m}^{*}(\beta_{c})$, at least for $\beta_{c}$ sufficiently large.
Finally we note that, to some extent, there is a similarity with the adsorption and collapse transition problem [VW96, VW98A, VW98B]. In particular, the fact that the density of mutual contacts is negligible below the mixing transition could suggest that some features of the boundaries between the phases are similar. For instance, we might guess that the phase boundary between the SE and the SC phases ($\theta$-point) does not depend on the value of $\beta_{m}$ when $\beta_{m}<\beta_{m}^{*}$.
@@ 3 Some rigorous results
exists in the hypercubic lattice for all $\beta_{c}<\infty$ [tesi1996interacting].
@@
and the limit exists for all $\beta_{c}<\infty$ [van1998collapsing].
@@
By the midpoint theorem [hardy1952inequalities] it follows that $\log Z_{2n}(\beta_{m},\beta_{c})$ is a convex function of $(\beta_{m},\beta_{c})$. Dividing by $2n$ and taking $n\to\infty$ gives the following theorem.
@@ 4 Monte Carlo method
Using a Markov chain Monte Carlo algorithm, we sampled the equilibrium conformations of the diblock lattice copolymer model from the Boltzmann distribution defined by equation (3). The elementary moves were a combination of pivot moves for self-avoiding polygons that ensure ergodicity [madras1990monte] and local Verdier-Stockmayer style moves [verdier1962monte] that increase the mobility of the Markov Chain, especially when the sampling is performed in the regions of large positive values of the interaction parameters $\beta_{m}$ and $\beta_{c}$ [tesi1996monte, tesi1996interacting].
Sampling was also improved by implementing the elementary moves using a Multiple Markov Chain algorithm [geyer1995annealing, tesi1996monte, tesi1996interacting] with parallel chains distributed along a sequence of parameters $(\beta_{c}{(j)},\beta_{m}{(j)})$ for $j=1,2,\ldots,M$. Along each parallel chain Metropolis sampling [metropolis1953monte] was implemented to sample from the Boltzmann distribution at a fixed pair of values $(\beta_{c}{(j)},\beta_{m}{(j)})$, and chains were swapped using the protocols of Multiple Markov Chain (MCC) sampling [Geyer:1991:CSS, tesi1996monte, tesi1996interacting].
@@
The stationary distribution of the MCC implementation is the product of the Boltzmann distributions along each chain (see reference [tesi1996monte, tesi1996interacting]). The main difficulty in the practical implementation of the MMC scheme is the choice of the sequence of parameter pairs $(\beta_{c}{(j)},)\beta_{m}{(j)}),j=1,2,\ldots,M$. This is because the contiguous pairs $(\beta_{c}{(j)},)\beta_{m}{(j)})$ and $(\beta_{c}{(j+1)},)\beta_{m}{(j+1)})$ should be sufficiently close to allow a non-negligible number of swaps between them to occur. In this study, we will estimate the metric and thermodynamic properties of the model by either fixing $\beta_{m}$ and varying $\beta_{c}$ or fixing $\beta_{c}$ and letting $\beta_{m}$ vary. More specifically, we choose a set of values of $\beta_{m}$ at fixed $\beta_{c}$ or a set of values of $\beta_{c}$ at fixed $\beta_{m}$ which samples the relevant region of the phase diagram and crosses the expected phase boundaries previously described.
@@ 5 Numerical results
In this section we explore in more detail the phase diagram illustrated in figure 2 using Monte Carlo simulations. We will in particular focus on the phase boundaries and the thermodynamical properties of the three phases.
@@ 5.1 The SE-SC phase boundary
Consider first the case that $\beta_{m}=0$. In this event the only nearest-neighbour self-attraction between monomers is when $\beta_{c}>0$. In the event that $\beta_{c}\leq 0$ the model is in a phase of segregated and expanded polygons. Increasing $\beta_{c}>0$ the model is expected to undergo a collapse transition at the $\theta$-point at a critical value $\beta_{c}^{*}>0$. In the case of cubic lattice polygons $\beta_{c}^{*}\approx 0.278$ [tesi1996interacting, bennett1998exact], and one expects this to be also the location of the $\theta$-point in our model when $\beta_{m}\leq 0$.
@@
Figure 6: (a) The locations of the peaks of $Var(k_{c})/2n$ for three values of the mutual attraction parameter $\beta_{m}=0.0,0.20,0.30$. (b) The corresponding locations of the crossings between the curve $\langle R_{g}^{2}\rangle/n^{2/3}$ vs $\beta_{c}$ at $n=400$ and the curves for $n=48,100,148,200$ and $300$. The three values of $\beta_{m}$ are the same as in (a). In both cases the extrapolation is against $1/\sqrt{n}$. The vertical bar at $1/\sqrt{n}=0$ refers to the estimate of $\beta_{c}^{*}$ reported for the standard $\theta$-point in [tesi1996interacting].
In figure 6(a) we plot the estimated locations of the maxima in the variances $Var(k_{c})/2n$. This is then extrapolated against $1/\sqrt{n}$ (at the $\theta$-point the finite size crossover exponent $\phi$ is expected to have its mean field value $1/2$). The crossings between the scaled mean square radius of gyration curves $\langle R_{g}^{2}\rangle/n^{2/3}$ for $n=400$ with the corresponding curves estimated for $n=48,100,148,200$ and $300$ are plotted in figure 6(b). The thick segment at $1/\sqrt{n}=0$ in the two graphs marks the estimated location of the cubic lattice $\theta$-point $\beta_{c}^{*}=0.2782\pm 0.007$ in reference [tesi1996interacting]. Observe that (i) the $n\to\infty$ extrapolations of the data in both graphs are very close to previous estimates of the cubic lattice $\theta$ transition; and (ii) the phase boundary between the SE and SC phases seems to be independent of the strength of the mutual attraction $\beta_{m}$, at least in the range $\beta_{m}\in[0,\beta_{m}^{*}]\approx[0,0.31]$.

### v2

@@ 1 Introduction
In this paper we investigate a model of two ring polymers that are required to be close together in space. The monomers in the two rings are different and this can lead to a mixed phase where the two rings interpenetrate [12]. In addition, either or both of the rings can collapse to form a compact ball. This model is connected to the behaviour of a diblock copolymer.
The synthesis of catenated ring polymers [24, 5], often in $AB$-diblock form [23] with one ring the $A$-block and the other ring the $B$-block, presents a novel situation with a phase diagram admitting several possible phases, including segregated, collapsed and mixed phases in a dilute solution, in addition to additional phases when the diblock catenane is in a melt or in confinement [5]. The linking of components in a catenane mix results in interlocking ring polymer sheets [18] resembling chainmail and held together by mechanically bonded rings due to the physical linking of the component ring polymers. Numerical simulations of catenated ring copolymers show that the rings are pulled together by catenation, giving rise to smaller radii of gyration (when compared to a single ring of the same length) and reduced distance between the centres of mass [27, 12]. Furthermore, it has been shown that ring-shaped polycatenanes have the capacity to retain a relevant quantity of twist, which significantly alters their metric and local properties [32].
@@
A cubic lattice model of two ring copolymers in a dilute solution that may be catenated consists of two interacting ring polymers (namely, lattice polygons) as illustrated in figure 1. The two polygons are held in proximity, and we examine metric and linking properties using Monte Carlo simulations. This approach is broadly similar to the study in reference [27], but with additional (repulsive or attractive) short-ranged forces between monomers in the two respective component polygons, which are also (independently) self-interacting if there are forces between monomers in each. This distinguishes the two polygons, so that, when catenated, they are a model of an $AB$-diblock ring polymer, with one ring the $A$-block and the other ring the $B$-block.
If the self-interacting forces between monomers in each component polygon are large, then they go through $\theta$-points to collapsed regimes, while, if the forces between monomers in each block (one in the $A$-block, and the other in the $B$-block) are large, then the two component polygons should interpenetrate each other, moving into a mixed regime. If either of the $A$- or $B$-blocks are collapsed, then they are in a collapsed or globular phase with increased self-entanglements as observed by the increased probability that they are non-trivial knots [30]. In the mixed regime the probability that the components form a catenane is increased [25, 26, 12]. We investigate this by examining the linking probability of the $A$- and $B$-blocks as a function of the interaction between the components.
We implement the model in figure 1 by considering two lattice polygons (or closed self-avoiding walks) of the same length on the simple cubic lattice. The two polygons are held together in the lattice by having at least one vertex in one polygon (this is the $A$-block) a unit distance from a vertex in the other polygon (namely the $B$-block). There are three energy terms in this model. If two monomers of type $A$ are unit distance apart and not connected by an edge of the polygon, then they form an $AA$-contact and they contribute an energy $\epsilon_{AA}$ to the total energy of the system. Similarly, two monomers of type $B$ a unit distance apart form a $BB$-contact and contribute an energy $\epsilon_{BB}$, and a pair of monomers of types $A$ and $B$ a unit distance apart form an $AB$-contact and contribute an energy $\epsilon_{AB}$. If $\epsilon_{AA}<0$ this will favour collapse of the $A$-ring into a globular phase, and similarly, $\epsilon_{BB}<0$ will favour collapse of the $B$-ring. If $\epsilon_{AB}<0$, this will favour intermingling of the two polygons and a transition to a mixed phase where we examine the linking probability of the two rings [12].
@@
If we set $\epsilon_{AA}=\epsilon_{BB}$, then both rings will collapse at the same temperature. In that case we expect at least three phases: A segregated-expanded phase where the rings are largely separated in space and each is expanded as in a good solvent, a segregated-collapsed phase where the two rings are largely separated but each ring has collapsed to a compact ball, and a mixed phase where the two rings interpenetrate to a significant extent. The phase diagram is illustrated in figure 2. We expect three phase boundaries in the phase diagram. The phase transition between the segregated-expanded and segregated-collapsed phases when $\epsilon_{AB}=0$ is relatively well understood. It is the collapse transition for a ring polymer.
[…]
We use several techniques to investigate this phase diagram. In section 3, we use rigorous arguments to establish the existence of a phase boundary as $\beta_{m}$ is increased, for all $\beta_{c}<\infty$. We write this phase boundary as $\beta_{m}=\beta_{m}^{*}(\beta_{c})$. In the same section, we show that, if a collapse transition exists as $\beta_{c}$ is increased with $\beta_{m}=0$ (say at $\beta_{c}=\beta_{c}^{*}$), then a transition occurs at the same value of $\beta_{c}$ for all $\beta_{m}<\beta_{m}(\beta_{c}^{*})$. It is thought that the transition at $\beta_{c}^{*}$ is second order and there are good numerical estimates of its location [3, 28]. The estimates are
@@
Our rigorous results leave many questions unanswered and we use Monte Carlo methods to obtain additional information. There are several quantities that one can examine to probe the details of the phase diagram. These include thermodynamic quantities such as the energy and heat capacity, and metric properties such as the radii of gyration of the rings (to detect collapse), and the mean distance between the centres of gravity of the rings (to detect mixing). In addition, we can probe the extent of intermingling or mutual entanglement by calculating the linking probability of the two rings. We proceed by sampling and collecting data on the model along lines crossing the phase boundaries in the phase diagram, using a multiple Markov Chain implementation [8, 29] of the Metropolis algorithm [21] using Verdier-Stockmayer [33] and pivot [19] elementary moves which can change the knot type of lattice polygons and linking complexity between lattice polygons.
@@ 2 The model
Mutual interactions between the two polygons are mutual contacts (pairs of vertices $(v_{A}\in A,v_{B}\in B)$ such that the distance between $v_{A}$ and $v_{B}$ is one lattice unit). As above, there is an interaction energy $\epsilon_{AB}$ for each mutual contact, and the total interaction energy from mutual contacts when a given configuration has $k_{m}$ mutual contacts is $E_{m}=\epsilon_{AB}k_{m}$. In addition, each block is also subject to an effective self-attraction due to the quality of the solvent. By defining a contact within block A as a pair of non-consecutive vertices that are a unit distance apart, we have $E_{A}=\epsilon_{AA}k_{A}$ where $\epsilon_{AA}<0$ is an effective attractive energy associated with each contact within block $A$ and $k_{A}$ is the number of such contacts. Similarly, for block $B$ we have $E_{B}=\epsilon_{BB}k_{B}$ with $\epsilon_{BB}<0$.
@@
In equation (2) the function $p_{2n}^{(2)}(k_{m},k_{A},k_{B})$ denotes the number of configurations of the two chemically distinct polygons with total length $2n$ having $k_{m}$ mutual contacts, and $k_{A}$, $k_{B}$ (self-)contacts within blocks $A$ and $B$ respectively. This general form accounts for the possibility of the two blocks having different affinities with the solvent. If we restrict to the simpler case of $\epsilon_{AA}=\epsilon_{BB}=\epsilon_{c}$ and define $k_{c}$ to be the total number of contacts, i.e. $k_{c}=k_{A}+k_{B}$, then the partition function (2) simplifies to
@@ 2.1 Preliminary considerations
Above a critical value $\beta_{m}^{*}$, the system is expected to be in a Mixed (M) phase where the two rings are significantly intermingled (the density of mutual contacts $\langle k_{m}\rangle/2n$ is now positive in the thermodynamic limit). The value of $\beta_{m}^{*}$ is $0.31$ for $\beta_{c}=0$  [12] but, for $\beta_{c}\neq 0$, we expect $\beta_{m}^{*}=\beta_{m}^{*}(\beta_{c})$, at least for $\beta_{c}$ sufficiently large.
Finally we note that, to some extent, there is a similarity with the adsorption and collapse transition problem [34, 36, 35]. In particular, the fact that the density of mutual contacts is negligible below the mixing transition could suggest that some features of the boundaries between the phases are similar. For instance, we might guess that the phase boundary between the SE and the SC phases ($\theta$-point) does not depend on the value of $\beta_{m}$ when $\beta_{m}<\beta_{m}^{*}$.
@@ 3 Some rigorous results
exists in the hypercubic lattice for all $\beta_{c}<\infty$ [28].
@@
and the limit exists for all $\beta_{c}<\infty$ [14].
@@
By the midpoint theorem [11] it follows that $\log Z_{2n}(\beta_{m},\beta_{c})$ is a convex function of $(\beta_{m},\beta_{c})$. Dividing by $2n$ and taking $n\to\infty$ gives the following theorem.
@@ 4 Monte Carlo method
Using a Markov chain Monte Carlo algorithm, we sampled the equilibrium conformations of the diblock lattice copolymer model from the Boltzmann distribution defined by equation (3). The elementary moves were a combination of pivot moves for self-avoiding polygons that ensure ergodicity [19] and local Verdier-Stockmayer style moves [33] that increase the mobility of the Markov Chain, especially when the sampling is performed in the regions of large positive values of the interaction parameters $\beta_{m}$ and $\beta_{c}$ [29, 28].
Sampling was also improved by implementing the elementary moves using a Multiple Markov Chain algorithm [8, 29, 28] with parallel chains distributed along a sequence of parameters $(\beta_{c}{(j)},\beta_{m}{(j)})$ for $j=1,2,\ldots,M$. Along each parallel chain Metropolis sampling [21] was implemented to sample from the Boltzmann distribution at a fixed pair of values $(\beta_{c}{(j)},\beta_{m}{(j)})$, and chains were swapped using the protocols of Multiple Markov Chain (MCC) sampling [9, 29, 28].
@@
The stationary distribution of the MCC implementation is the product of the Boltzmann distributions along each chain (see reference [29, 28]). The main difficulty in the practical implementation of the MMC scheme is the choice of the sequence of parameter pairs $(\beta_{c}{(j)},)\beta_{m}{(j)}),j=1,2,\ldots,M$. This is because the contiguous pairs $(\beta_{c}{(j)},)\beta_{m}{(j)})$ and $(\beta_{c}{(j+1)},)\beta_{m}{(j+1)})$ should be sufficiently close to allow a non-negligible number of swaps between them to occur. In this study, we will estimate the metric and thermodynamic properties of the model by either fixing $\beta_{m}$ and varying $\beta_{c}$ or fixing $\beta_{c}$ and letting $\beta_{m}$ vary. More specifically, we choose a set of values of $\beta_{m}$ at fixed $\beta_{c}$ or a set of values of $\beta_{c}$ at fixed $\beta_{m}$ which samples the relevant region of the phase diagram and crosses the expected phase boundaries previously described.
@@ 5 Numerical results
In this section we explore in more detail the phase diagram illustrated in figure 2 using Monte Carlo simulations. We will in particular focus on the phase boundaries and the thermodynamic properties of the three phases.
@@ 5.1 The SE-SC phase boundary
Consider first the case that $\beta_{m}=0$. In this event the only nearest-neighbour self-attraction between monomers is when $\beta_{c}>0$. In the event that $\beta_{c}\leq 0$ the model is in a phase of segregated and expanded polygons. Increasing $\beta_{c}>0$ the model is expected to undergo a collapse transition at the $\theta$-point at a critical value $\beta_{c}^{*}>0$. In the case of cubic lattice polygons $\beta_{c}^{*}\approx 0.278$ [28, 3], and one expects this to be also the location of the $\theta$-point in our model when $\beta_{m}\leq 0$.
@@
Figure 6: (a) The locations of the peaks of $Var(k_{c})/2n$ for three values of the mutual attraction parameter $\beta_{m}=0.0,0.20,0.30$. (b) The corresponding locations of the crossings between the curve $\langle R_{g}^{2}\rangle/n^{2/3}$ vs $\beta_{c}$ at $n=400$ and the curves for $n=48,100,148,200$ and $300$. The three values of $\beta_{m}$ are the same as in (a). In both cases the extrapolation is against $1/\sqrt{n}$. The vertical bar at $1/\sqrt{n}=0$ refers to the estimate of $\beta_{c}^{*}$ reported for the standard $\theta$-point in [28].
In figure 6(a) we plot the estimated locations of the maxima in the variances $Var(k_{c})/2n$. This is then extrapolated against $1/\sqrt{n}$ (at the $\theta$-point the finite size crossover exponent $\phi$ is expected to have its mean field value $1/2$). The crossings between the scaled mean square radius of gyration curves $\langle R_{g}^{2}\rangle/n^{2/3}$ for $n=400$ with the corresponding curves estimated for $n=48,100,148,200$ and $300$ are plotted in figure 6(b). The thick segment at $1/\sqrt{n}=0$ in the two graphs marks the estimated location of the cubic lattice $\theta$-point $\beta_{c}^{*}=0.2782\pm 0.007$ in reference [28]. Observe that (i) the $n\to\infty$ extrapolations of the data in both graphs are very close to previous estimates of the cubic lattice $\theta$ transition; and (ii) the phase boundary between the SE and SC phases seems to be independent of the strength of the mutual attraction $\beta_{m}$, at least in the range $\beta_{m}\in[0,\beta_{m}^{*}]\approx[0,0.31]$.